Santha is 10 years old. His little sister, Shurti, is only two years old. How many times older than Shruti is Santha?

A. 10 = 2(X)
B. 10 = 2 + 2(X)
C. None of the above

If your answer choice is C., then write a number sentence to represent the story correctly.
________________________________________________________________

I think the answer is B, because Shurti is 2 and Santha is 8 years older, so that would be 2 + 2(4) = 10

Answer by wilft1(217)

the answer would be A
10/2 = 5, so 10 = 2(x) would be the correct equation
lets solve b though to help understand...
10 = 2 + 2(x)
subtract 2 from both sides
8 = 2(x)
divide both sides by 2
4 = x
2 times 4 = 8
santha is 10 years old, not 8 years old. if santha was 8 years old then B would work.
my email is wilft1@yahoo.com i dont know what you mean by a bar model